Sun Dance Gardens Inc. develops and produces spraying equipment for lawn maintenance and industrial uses. On June 3 of the current year, Sun Dance Gardens Inc. reacquired 14,900 shares of its common stock at $32 per share.

a. What is the balance of Treasury Stock on December 31 of the current year?

b. How will the balance in Treasury Stock be reported on the balance sheet?

c. Assume that Sun Dance Gardens sold 9,400 shares of its treasury stock at $34 on November 2. What accounts would be affected by the sale of the treasury stock?

Paid-In Capital from Treasury Stock by $

Cash by $

Treasury Stock by $
Answer #1

a) Balance in treasury Stock account = 14900*32 = 476800

b) Deducted from total Stockholder's equity

c) Effect on account

Account Effect Amount
Paid in capital from treasury stock Increase 9400*2 = 18800
Cash Increase 9400*34 = 319600
Treasury stock Decrease 9400*32 = 300800